For any positive integer n, ((n+1)!/n!)-n =
The answer is 1 Please explain....

Answer by rothauserc(4718)

((n+1)!/n!)-n =
work the terms inside the parenthesis first
note that (n+1)! = (n+1)*n!, so
(n+1)*(n)! / n! = n+1 because the n! terms cancel
we are left with
n+1 - n = 1